OpenStudy (anonymous):

The slope of the line containing the points (-5, 3) and (-2, 1) is

OpenStudy (luigi0210):

Use the slope formula: \(\Large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\)

OpenStudy (anonymous):

tried and its not coming out right for me I have 3 answeres it can be and not getting any -2/3, 2/3 -3/2

OpenStudy (anonymous):

Let me do it

OpenStudy (anonymous):

2/-7

OpenStudy (anonymous):

Wait no nevermind

OpenStudy (anonymous):

3-1/5-(-2) = 2/7

OpenStudy (anonymous):

these are the 3 I can choose from -3/2 2/3 -2/3

OpenStudy (anonymous):

2/3

OpenStudy (luigi0210):

Did you plug the points into the right places? \[\Large m=\frac{1-3}{-2-(-5)}\] No giving direct answers please~

OpenStudy (anonymous):

1-3 and -2-(-5) so -2/3

OpenStudy (anonymous):

yes I did and am getting 35/2

OpenStudy (luigi0210):

Where did that come from?

OpenStudy (anonymous):

I got it thank you

OpenStudy (anonymous):

Me and luigi gave you the answer. Solve luigis equation